is this the correct way of doing quadratic equation
x^2+x-6=0
(x+2)(x-3)
x+2=0 x-3=0
x=2 x=3

Answer by Alan3354(69443)

is this the correct way of doing quadratic equation
x^2+x-6=0
(x+2)(x-3)
x+2=0 x-3=0
x=2 x=3
---------------
It's a correct way, when you can factor the polynomial.
-----
x^2+x-6=0
(x+2)(x-3) ***** = 0
x+2=0 x-3=0
x=2 **** x = -2
x=3